1.Suppose that 2 years ago you bought an old record player at a yard sale for $1. You saw today on E-bay that the same record player is selling for $53. If you were to sell the record play at that price today, what would be the implied return percentage? (Convert to a percent. Round to 2 decimal places.

2.in 1998, the average price of a gallon of gas was $1.07. Today, the average price of a gallon of gas is $2.71. At what annual rate has a gallon of gas increased over the last 20 years? (Answer as a percent. Enter only numbers and decimals in your response. Round to 2 decimal places.)

3.An investment will pay you $4,903 in 4 years if you pay $1,729 today. What is the implied rate of return? (Convert to a decimale. Round to 2 decimal places.)

please answer all the questions. i will be very appropriate :)
Answer #1

Answer 1.

Price, 2 years ago = $1.00
Current price = $53.00

Implied return percentage = (Current price / Price, 2 years ago)^(1/2) - 1
Implied return percentage = ($53.00 / $1.00)^(1/2) - 1
Implied return percentage = 53^(1/2) - 1
Implied return percentage = 7.2801 - 1
Implied return percentage = 6.2801 or 628.01%

Answer 2.

Price, 20 years ago = $1.07
Current price = $2.71

Implied return percentage = (Current price / Price, 20 years ago)^(1/20) - 1
Implied return percentage = ($2.71 / $1.07)^(1/20) - 1
Implied return percentage = 2.532710^(1/20) - 1
Implied return percentage = 1.0476 - 1
Implied return percentage = 0.0476 or 4.76%

Answer 3.

Present value = $1,729
Future value = $4,903
Time period = 4 years

Implied return percentage = (Future value / Present value)^(1/Time period) - 1
Implied return percentage = ($4,903 / $1,729)^(1/4) - 1
Implied return percentage = 2.835743^(1/4) - 1
Implied return percentage = 1.2977 - 1
Implied return percentage = 0.2977 or 29.77%